# IMPACT: Behavioral Intention-aware Multimodal Trajectory Prediction with Adaptive Context Trimming

### v1

@@ 1 Introduction
Figure 1: (a) illustrates the traditional predictor’s context input, while (b) is our integrated approach jointly predicting behavioral intentions, trajectories, and vectorized occupancy. In our approach, the decoder stage is fed only with influential agents and relevant map elements.
@@
Currently, most trajectory prediction models use attention mechanism for querying information from agents and map features. However, indiscriminately attending to all agents and map elements introduces unnecessary complexity and potential causal confusion, as many elements have no direct relevance to the target agent, ultimately hindering correct decoding of the target agent’s future trajectories. To address this, some recent approaches introduce an additional local context-aware refinement module after the decoder stage [43, 6], but this further complicates the prediction pipeline. Others adopt rule-based heuristic pruning of input lanes [17] or dynamically collect map elements based on the previous layer prediction results [26, 27, 19, 29]. However, these methods may struggle in complex scenarios or suffer from error propagation. BETOP [20] is the first paper explicitly modelling agent interactions through braid theory, but its topological selection algorithm demonstrates critical accuracy deficiencies (see appendix). Thus, an efficient approach to selectively focus on truly critical features (both agents and maps) remains an open challenge. Furthermore, the lack of ground-truth intention labels in current mainstream datasets (e.g., Waymo [38], Argoverse [1, 2]) is limiting the development of accurate behavioral intention-aware models.
To holistically address the aforementioned challenges, we propose our approach, IMPACT. Specifically, we introduce an agent-wise multimodal behavioral intention module before the trajectory decoder, while sharing the same scenario information extraction module (i.e., the same encoder) used by the trajectory predictor. This design reduces computational overhead and mitigates information loss. The agent-wise behavioral intention module outputs a one-hot intention vector for each vehicle, where each element in the vector represents the probability of a specific intention relative to the ego vehicle (e.g., overtaking, yielding). Based on these probability vectors, we apply a utility function to assign an “interaction score” to each vehicle, selecting only those with high interaction probabilities as inputs to the decoder. In parallel, we propose a vectorized occupancy prediction module. This module predicts the probability that each vectorized map element will be occupied by the ego vehicle in the future. Similar to the agent selection process, we only feed the map polylines with high occupancy probabilities into the decoder. By integrating both modules in this manner, our method effectively reduces redundant interactions while preserving crucial information for accurate and robust trajectory prediction. We also propose an automatic labeling algorithm to generate high-quality ground truth labels for behavioral intentions on the Waymo and Argoverse 1&2 dataset.
@@ 2.1 Motion Prediction
However, current methods tend to emphasize the precise generation of multimodal trajectory predictions at the expense of behavioral intention prediction[8] (BIP). This narrow focus often overlooks the latent decision-making processes that drive observable maneuvers, resulting in models that struggle to interpret complex social interactions or anticipate nuanced changes in driving behavior. In real-world scenarios, the inability to infer intentions such as overtake or yield decisions can lead to suboptimal planning and reduced overall robustness. The most recent work, BeTOP [20], attempts to explicitly model behavioral interactions via braid theory by splitting agents into “interactive” and “non-interactive” based on whether their future trajectories form an intertwined braid. However, in most scenarios (see Appendix), the generated braids appear unrealistic because the approach only considers lateral and temporal dimensions. Moreover, relying on a binary behavioral label oversimplifies complex traffic scenarios, where interactions can be far more nuanced.
@@ 3.2 Input Representation
Following the approaches of RMP-YOLO[29] and MacFormer[9], we further incorporate the historical relative movement between the target agent and the map polylines to capture dynamic subtle interdependencies. This historical movement is denoted by =$\mathcal{R}=\{r_{1},r_{2},\dots,r_{N_{l}}\}\in\mathbb{R}^{N_{l}\times T_{p}\times F_{3}},$ where $F_{3}$ is the feature size associated with the relative movement($(\Delta x,\,\Delta y,\,\cos\Delta\theta,\,\sin\Delta\theta)$).
@@ 3.3.3 Multimodal Behavioral Intention Prediction
Each vector element $h$ represents a probability distribution over four intention categories: yielding, overtaking, ignored, and nearby. To make the decoding process more focused, we first compute an overall interaction score from the predicted distribution $\hat{\mathrm{H}}$ using a utility function $\psi$, producing $\psi(\hat{\mathrm{H}^{i}})\in\mathbb{R}^{K\times N_{a}}$. We then select the top-$m$ highest-scoring agents for downstream trajectory decoding. Therefore, for each modality, we choose $m$ most revelant agents $A_{5}\in\mathbb{R}^{m\times D}$, where $m\ll N_{a}$. This filtering step refines the decoder’s input, concentrating on the most influential interactions while improving prediction accuracy. The ground-truth label of behavioral intention $\mathrm{H}^{*}$ is derived from an auto-labeled data preprocessing process (see Appendix).
@@ 3.3.4 Multimodal Vectorized Occupancy Prediction
This vectorized approach ensures both efficiency and scalability while maintaining alignment with the overall vectorized representation of the scene. Among the $N_{l}$ polylines’s multimodal occupancy probabilities $\varphi(\hat{\mathrm{O}^{i}})\in\mathbb{R}^{K\times N_{l}}$, we select the top-$n$ with the highest predicted probabilities in each modality to form $L_{5}\in\mathbb{R}^{n\times D}$, where $n\ll N_{l}$. These top-ranked polylines serve as focused inputs for the subsequent trajectory decoder. The ground-truth occupancy label $\mathrm{O}^{*}$ is also derived from an auto-labeled data preprocessing process (see Appendix).
@@ 3.4 Training Loss
where $\lambda_{\mathrm{1}},\lambda_{\mathrm{2}},\lambda_{\mathrm{3}}$, and $\lambda_{\mathrm{4}}$ are weighting factors balancing the contributions of behavioral intention prediction, vectorized occupancy prediction, trajectory prediction, and mode classification, respectively. Specifically, $\mathcal{L}_{\mathrm{Int}}$ is calculated using the multi-class Focal Loss, $\mathcal{L}_{\mathrm{Occ}}$ is based on the binary Focal Loss, $\mathcal{L}_{\mathrm{Traj}}$ is derived from the GMM loss, and $\mathcal{L}_{\mathrm{Score}}$ is computed with Binary Cross-Entropy. During training, the winner-take-all strategy is applied for $\mathcal{L}_{\mathrm{Int}}$, $\mathcal{L}_{\mathrm{Occ}}$, and $\mathcal{L}_{\mathrm{Traj}}$, ensuring that only the modality closest to the ground-truth trajectory is used to compute these losses. Please check appendix for more details.
@@ 4.1 Experimental Setup
We employ AdamW optimizer [21] for training, conducting experiments on a cluster of 8 NVIDIA A800 GPUs with a total batch size of 80. The learning rate is initialized as $1\times 10^{-4}$ and begins step decay starting at epoch 22, halving every two epochs. The model undergoes 30 epochs. More details can be referred to Appendix.
@@ 4.2 Leaderboard Performance
Figure 4 illustrates the detailed qualitative results for both marginal and joint predictions using our proposed method. LABEL:fig:figure-sampling-space shows our qualitative marginal and joint prediction results. For additional qualitative results, please refer to Appendix.
Marginal Prediction Performance.
Joint Prediction Performance.
4.3 Behavioral Intention Prediction Performance
Table 1: Multimodal Behavioral Intention Prediction Performance.
Top-1 Top-6
Class Precision Recall F1-Score Accuracy GT Data Ratio (%)
Ignored 0.99 0.97 0.98 0.99 89.12
Nearby 0.71 0.89 0.79 0.97 3.93
Overtaking 0.82 0.89 0.85 0.97 2.59
Yielding 0.86 0.96 0.90 0.97 4.32
All 0.97 0.97 0.97 0.987 100.00
4.4 Vectorized Occupancy Prediction Performance
Class Precision Recall F1-Score GT Data Ratio (%)
Occupied 0.933 0.775 0.847 3.96
Unoccupied 0.991 0.998 0.994 96.04
All 0.989 100.00
Table 2 presents precision, recall, and F1-score for both classes, along with their data ratios in the validation set. The occupied class constitutes only 3.96% of samples, highlighting a strong imbalance. Despite this, the model achieves a high overall accuracy of 0.989. For the occupied class, precision (0.933) and recall (0.775) yield an F1-score of 0.847, indicating rare false positives—crucial for pruning irrelevant map polylines in trajectory prediction. The dominant unoccupied class achieves near-perfect precision (0.991), recall (0.998), and F1-score (0.994). These results demonstrate that our vectorized occupancy prediction effectively identifies critical map segments while minimizing false classification. More qualitative results are in Appendix.
@@ 4.5 Ablation Study
On average, the model input contains 43.85 agents and 749.52 map polylines, meaning this selection reduces scene complexity by 45.2% and 74%, respectively, while preserving essential contextual cues. Beyond this threshold, redundant elements introduce noise, degrading generalization and increasing prediction uncertainty. This highlights the importance of selective context pruning to enhance trajectory forecasting performance.
@@
We compare various agent-pruning and map-pruning strategies in Table 5 using the same encoder architecture. The first row (All + w/Dynamic) corresponds to a baseline that queries all agents and dynamically collect polylines, resulting in a model with 69.469 M parameters and a reported inference time of 14.67 ms under a simplified setting. By contrast, our final approach (the last row in Table 5) combines YOIN-based agent pruning with vectorized occupancy-based pruning. This design lowers the parameter count to 48.944 M while reducing the average inference time of BeTOP[20] from 23.92 to 18.27 ms per scenario. The key to these efficiency gains lies in pruning irrelevant scene information—both agents and polylines—based on their semantic contributions, which in turn eliminates redundant computations. Notably, the approach retains or improves accuracy metrics such as Soft mAP, mAP, and minADE, underscoring that targeted context selection effectively balances computational efficiency with predictive performance.
Table 6: Performance gains from integrating proposed modules with existing approaches (Argoverse 1 & 2). bFDE6 is the main metric.

### v2

@@ 1 Introduction
Figure 1: (a) illustrates the traditional context input, while (b) is our integrated approach jointly predicting behavioral intentions, trajectories, and vectorized occupancy. In our approach, the decoder stage is fed only with influential agents and relevant map elements.
@@
Currently, most trajectory prediction models use attention mechanism for querying information from agents and map features. However, indiscriminately attending to all agents and map elements introduces unnecessary complexity and potential causal confusion, as many elements have no direct relevance to the target agent, ultimately hindering correct decoding of the target agent’s future trajectories. To address this, some recent approaches introduce an additional local context-aware refinement module after the decoder stage [45, 6], but this further complicates the prediction pipeline. Others adopt rule-based heuristic pruning of input lanes [17] or dynamically collect map elements based on the previous layer prediction results [27, 28, 19, 30]. However, these methods may struggle in complex scenarios or suffer from error propagation. BETOP [21] is the first paper explicitly modelling agent interactions through braid theory, but its topological selection algorithm demonstrates critical accuracy deficiencies (see appendix Figure 9). Thus, an efficient approach to selectively focus on truly critical features (both agents and maps) remains an open challenge. Furthermore, the lack of ground-truth intention labels in current mainstream datasets (e.g., Waymo [39], Argoverse [1, 2]) is limiting the development of accurate behavioral intention-aware models.
To holistically address the aforementioned challenges, we propose our approach, IMPACT. Specifically, we introduce an agent-wise multimodal behavioral intention module before the trajectory decoder, while sharing the same scenario information extraction module (i.e., the same encoder) used by the trajectory predictor. This design reduces computational overhead and mitigates information loss. The agent-wise behavioral intention module outputs a one-hot intention vector for each vehicle, where each element in the vector represents the probability of a specific intention relative to the target vehicle (e.g., overtaking, yielding). Based on these probability vectors, we apply a utility function to assign an “interaction score” to each vehicle, selecting only those with high interaction probabilities as inputs to the decoder. In parallel, we propose a vectorized occupancy prediction module. This module predicts the probability that each vectorized map element will be occupied by the target vehicle in the future. Similar to the agent selection process, we only feed the map polylines with high occupancy probabilities into the decoder. By integrating both modules in this manner, our method effectively reduces redundant interactions while preserving crucial information for accurate and robust trajectory prediction. We also propose an automatic labeling algorithm to generate high-quality ground truth labels for behavioral intentions on the Waymo and Argoverse 1&2 dataset.
@@ 2.1 Motion Prediction
However, current methods tend to emphasize the precise generation of multimodal trajectory predictions at the expense of behavioral intention prediction[8] (BIP). This narrow focus often overlooks the latent decision-making processes that drive observable maneuvers, resulting in models that struggle to interpret complex social interactions or anticipate nuanced changes in driving behavior. In real-world scenarios, the inability to infer intentions such as overtake or yield decisions can lead to suboptimal planning and reduced overall robustness. The most recent work, BeTOP [21], attempts to explicitly model behavioral interactions via braid theory by splitting agents into “interactive” and “non-interactive” based on whether their future trajectories form an intertwined braid. However, in most scenarios (see Appendix Figure 9), the generated braids appear unrealistic because the approach only considers lateral and temporal dimensions. Moreover, relying on a binary behavioral label oversimplifies complex traffic scenarios, where interactions can be far more nuanced.
@@ 3.2 Input Representation
Following the approaches of ControlMTR[31], RMP-YOLO[30], and MacFormer[9], we further incorporate the historical relative movement between the target agent and the map polylines to capture dynamic subtle interdependencies. This historical movement is denoted by =$\mathcal{R}=\{r_{1},r_{2},\dots,r_{N_{l}}\}\in\mathbb{R}^{N_{l}\times T_{p}\times F_{3}},$ where $F_{3}$ is the feature size associated with the relative movement($(\Delta x,\,\Delta y,\,\cos\Delta\theta,\,\sin\Delta\theta)$).
@@ 3.3.3 Multimodal Behavioral Intention Prediction
Each vector element $h$ represents a probability distribution over four intention categories: yielding, overtaking, ignored, and nearby. To make the decoding process more focused, we first compute an overall interaction score from the predicted distribution $\hat{\mathrm{H}}$ using a utility function $\psi$, producing $\psi(\hat{\mathrm{H}^{i}})\in\mathbb{R}^{K\times N_{a}}$. We then select the top-$m$ highest-scoring agents for downstream trajectory decoding. Therefore, for each modality, we choose $m$ most revelant agents $A_{5}\in\mathbb{R}^{m\times D}$, where $m\ll N_{a}$. This filtering step refines the decoder’s input, concentrating on the most influential interactions while improving prediction accuracy. The ground-truth label of behavioral intention $\mathrm{H}^{*}$ is derived from an auto-labeled data preprocessing process (see Appendix Algorithm 2).
@@ 3.3.4 Multimodal Vectorized Occupancy Prediction
This vectorized approach ensures both efficiency and scalability while maintaining alignment with the overall vectorized representation of the scene. Among the $N_{l}$ polylines’s multimodal occupancy probabilities $\varphi(\hat{\mathrm{O}^{i}})\in\mathbb{R}^{K\times N_{l}}$, we select the top-$n$ with the highest predicted probabilities in each modality to form $L_{5}\in\mathbb{R}^{n\times D}$, where $n\ll N_{l}$. These top-ranked polylines serve as focused inputs for the subsequent trajectory decoder. The ground-truth occupancy label $\mathrm{O}^{*}$ is also derived from an auto-labeled data preprocessing process (see Appendix Algorithm 1).
@@ 3.4 Training Loss
where $\lambda_{\mathrm{1}},\lambda_{\mathrm{2}},\lambda_{\mathrm{3}}$, and $\lambda_{\mathrm{4}}$ are weighting factors balancing the contributions of behavioral intention prediction, vectorized occupancy prediction, trajectory prediction, and mode classification, respectively. Specifically, $\mathcal{L}_{\mathrm{Int}}$ is calculated using the multi-class Focal Loss, $\mathcal{L}_{\mathrm{Occ}}$ is based on the binary Focal Loss, $\mathcal{L}_{\mathrm{Traj}}$ is derived from the GMM loss, and $\mathcal{L}_{\mathrm{Score}}$ is computed with Binary Cross-Entropy. During training, the winner-take-all strategy is applied for $\mathcal{L}_{\mathrm{Int}}$, $\mathcal{L}_{\mathrm{Occ}}$, and $\mathcal{L}_{\mathrm{Traj}}$, ensuring that only the modality closest to the ground-truth trajectory is used to compute these losses. Please check appendix (section 2) for more details.
@@ 4.1 Experimental Setup
We employ AdamW optimizer [22] for training, conducting experiments on a cluster of 8 NVIDIA A800 GPUs with a total batch size of 80. The learning rate is initialized as $1\times 10^{-4}$ and begins step decay starting at epoch 22, halving every two epochs. The model undergoes 30 epochs. More details can be referred to Appendix section 4.
@@ 4.2 Leaderboard Performance
Figure 5 illustrates the detailed qualitative results for both marginal and joint predictions using our proposed method. For additional qualitative results, please refer to section 7 in Appendix.
@@ 4.3 Ablation Study
Marginal Prediction Performance.
Joint Prediction Performance.
Table 4: Multimodal Behavioral Intention Prediction Performance.
Top-1 Top-6
Class Precision Recall F1-Score Accuracy GT Data Ratio (%)
Ignored 0.99 0.97 0.98 0.99 89.12
Nearby 0.71 0.89 0.79 0.97 3.93
Overtaking 0.82 0.89 0.85 0.97 2.59
Yielding 0.86 0.96 0.90 0.97 4.32
All 0.97 0.97 0.97 0.987 100.00
4.4 Behavioral Intention Prediction Performance
4.5 Vectorized Occupancy Prediction Performance
Class Precision Recall F1-Score GT Data Ratio (%)
Occupied 0.933 0.775 0.847 3.96
Unoccupied 0.991 0.998 0.994 96.04
All 0.989 100.00
Table 5 presents precision, recall, and F1-score for both classes, along with their data ratios in the validation set. The occupied class constitutes only 3.96% of samples, highlighting a strong imbalance. Despite this, the model achieves a high overall accuracy of 0.989. For the occupied class, precision (0.933) and recall (0.775) yield an F1-score of 0.847, indicating rare false positives—crucial for pruning irrelevant map polylines in trajectory prediction. The dominant unoccupied class achieves near-perfect precision (0.991), recall (0.998), and F1-score (0.994). These results demonstrate that our vectorized occupancy prediction effectively identifies critical map segments while minimizing false classification. More qualitative results (like Figure 4) are in Appendix section 7.
@@ 4.5 Vectorized Occupancy Prediction Performance
As shown in Figure 6, selecting 24 agents and 192 polylines achieves the optimal balance between interaction diversity and trajectory accuracy. On average, the model input contains 43.85 agents and 749.52 map polylines, meaning this selection reduces scene complexity by 45.2% and 74%, respectively, while preserving essential contextual cues. Beyond this threshold, redundant elements introduce noise, degrading generalization and increasing prediction uncertainty. This highlights the importance of selective context pruning to enhance trajectory forecasting performance.
@@
We compare various agent-pruning and map-pruning strategies in Table 3 using the same encoder architecture. The first row (All + w/Dynamic) corresponds to a baseline that queries all agents and dynamically collect polylines, resulting in a model with 69.469 M parameters and a reported inference time of 14.67 ms under a simplified setting. By contrast, our final approach (the last row in Table 3) combines YOIN-based agent pruning with vectorized occupancy-based pruning. This design lowers the parameter count to 48.944 M while reducing the average inference time of BeTOP[21] from 23.92 to 18.27 ms per scenario. The key to these efficiency gains lies in pruning irrelevant scene information—both agents and polylines—based on their semantic contributions, which in turn eliminates redundant computations.
Table 6: Performance gains from integrating proposed modules with existing approaches(Argoverse1&2). bFDE6 is main metric.